# Diffusion-SDPO: Safeguarded Direct Preference Optimization for Diffusion Models

### v1

@@ 6 Conclusions and Limitations
In this paper, we presented Diffusion-SDPO, a safeguarded preference optimization scheme that stabilizes DPO-style diffusion finetuning by preserving the preferred branch while improving preference matching. The method scales the loser gradient by its alignment with the winner and guarantees, to first order, that the winner’s reconstruction loss does not increase. Across SD 1.5, SDXL, and Ovis-U1, our method yields consistent improvements on automated preference, aesthetic, and prompt-alignment metrics with negligible computational overhead, while remaining model-agnostic, straightforward to implement, and applicable to multiple DPO variants. However, the guarantee holds only at first order, so strong curvature in the loss landscape degrades it. Future work includes second-order or trust-region safeguards, automatic or data-driven tuning of $\mu$, extensions to autoregressive preference settings.
Reproducibility Statement
All datasets and base models used in this work are publicly available except for a small mixed preference set that we curate from publicly accessible sources and use only for the Ovis-U1 experiments. Sec. 5.1 details the models, datasets, as well as the complete training and evaluation protocols. Sec. 5.3 reports ablations and sensitivity analyses. Core results on SD 1.5 and SDXL rely solely on public data and base models. The unified Ovis-U1 model exhibits empirical trends consistent with results from SD 1.5 and SDXL. With the documented settings, hyperparameters, and evaluation procedures, independent researchers can reproduce the SD 1.5 and SDXL results using public resources.
Ethics Statement
We adhere to the ICLR Code of Ethics. Our experiments rely primarily on public datasets and model releases. For one unified model variant, we curated a small mixed preference set from publicly accessible sources and applied basic filtering for unsafe content where practicable. No human-subjects studies were conducted and, to our knowledge, no personally identifiable information was processed, so institutional review was not required.
Appendix
Appendix A Justification of the near–isometry
In this section, we provide a detailed justification of the near-isometry property $J^{w}{J^{l}}^{\top}\approx I$. We write the model output at diffusion step $t$ as $o=\epsilon_{\theta}(x_{t},c,t)$. For the winner and loser branches we denote
$o^{w}=\epsilon_{\theta}(x_{t}^{w},c,t),\qquad o^{l}=\epsilon_{\theta}(x_{t}^{l},c,t),\qquad J^{w}=\frac{\partial o^{w}}{\partial\theta},\qquad J^{l}=\frac{\partial o^{l}}{\partial\theta}.$ (18)
By first–order linearization at the same parameter $\theta$,
$\nabla_{\theta}\mathcal{L}^{w}={J^{w}}^{\top}\nabla_{o}\mathcal{L}^{w},\qquad\nabla_{\theta}\mathcal{L}^{l}={J^{l}}^{\top}\nabla_{o}\mathcal{L}^{l}.$ (19)
Hence the parameter–space inner product factorizes as
@@ Appendix A Justification of the near–isometry
$\nabla_{\theta}{\mathcal{L}^{w}}^{\top}\nabla_{\theta}\mathcal{L}^{l}=\nabla_{o}{\mathcal{L}^{w}}^{\top}\,\underbrace{(J^{w}{J^{l}}^{\top})}_{\text{geometric term}}\,\nabla_{o}\mathcal{L}^{l}.$ (20)
We claim that in our setting $J^{w}{J^{l}}^{\top}\approx I$, which makes the parameter–space geometry well approximated by the Euclidean geometry in output space. Below we state the assumptions and a quantitative bound that together justify this approximation.
Assumption A (branch proximity).
The two branches are evaluated under the same condition $(c,t)$. For a fixed $t$, the noised latents $x_{t}^{w}$ and $x_{t}^{l}$ lie in a small neighborhood in latent space. The UNet unet is Lipschitz in its input around that neighborhood, therefore the Jacobians are close:
$\|J^{l}-J^{w}\|\leq\varepsilon_{\text{br}},\qquad\varepsilon_{\text{br}}\ll 1.$ (21)
This follows from sharing the text condition, the timestep, and the architecture, which produce similar intermediate activations and hence similar sensitivities to parameters.
Assumption B (single–branch near–isometry).
Each branch is locally well conditioned. Empirically, modern diffusion UNets use normalization layers, residual connections, and near unit–gain initializations. Besides, they are trained to predict isotropic Gaussian noise at each step. These design choices constrain the singular values of $J$ to stay close to one. We model this as
$\|J^{w}{J^{w}}^{\top}-I\|\leq\varepsilon_{\text{iso}},\qquad\|J^{l}{J^{l}}^{\top}-I\|\leq\varepsilon_{\text{iso}},$ (22)
with $\varepsilon_{\text{iso}}\ll 1$. In words, the map $\theta\mapsto o$ is locally near–isometric.
A perturbation bound.
Let $E_{\text{br}}:=J^{l}-J^{w}$ and $E_{\text{iso}}:=J^{w}{J^{w}}^{\top}-I$. Then
$J^{w}{J^{l}}^{\top}-I=J^{w}(J^{w}+E_{\text{br}})^{\top}-I=\underbrace{(J^{w}{J^{w}}^{\top}-I)}_{E_{\text{iso}}}+J^{w}E_{\text{br}}^{\top}.$ (23)
Taking operator norms and using $\|AB\|\leq\|A\|\,\|B\|$ gives
$\|J^{w}{J^{l}}^{\top}-I\|\;\leq\;\|E_{\text{iso}}\|+\|J^{w}\|\,\|E_{\text{br}}\|\;\leq\;\varepsilon_{\text{iso}}+(1+\varepsilon_{\text{gain}})\,\varepsilon_{\text{br}},$ (24)
where $\varepsilon_{\text{gain}}\geq 0$ is defined by $\|J^{w}\|\leq 1+\varepsilon_{\text{gain}}$. If both branches are individually near–isometric and their Jacobians are close, then the cross term $J^{w}{J^{l}}^{\top}$ is close to the identity.
Consequence for gradient geometry.
Combining Eq. 20 and 24 yields
$\Big\|\nabla_{\theta}{\mathcal{L}^{w}}^{\top}\nabla_{\theta}\mathcal{L}^{l}-\nabla_{o}{\mathcal{L}^{w}}^{\top}\nabla_{o}\mathcal{L}^{l}\Big\|\;\leq\;\|\nabla_{o}\mathcal{L}^{w}\|\,\|\nabla_{o}\mathcal{L}^{l}\|\cdot\|J^{w}{J^{l}}^{\top}-I\|,$ (25)
so the inner product between parameter–space gradients is close to the inner product between output–space gradients. Intuitively, in the small neighborhood induced by the shared $(c,t)$, the network behaves almost angle– and length–preserving along the two branches, which justifies replacing the parameter–space inner product by the Euclidean dot product in output space.
Why Assumption B is plausible for diffusion UNets.
At a fixed $t$, the training objective is a mean–squared error between the model output and an isotropic Gaussian target. This induces a whitened curvature in output space. Together with normalization layers and residual pathways, which are known to promote dynamical isometry by controlling singular values batchnorm; dynamic, the per–step Jacobian spectrum concentrates around one. This effect is stable under small finetuning steps and within the local region defined by $x_{t}^{w}$ and $x_{t}^{l}$.
Empirical diagnostics.
We validate the approximation by comparing the cosine between parameter space gradients $\rho_{\theta}=\cos\!\big(\nabla_{\theta}\mathcal{L}^{w},\nabla_{\theta}\mathcal{L}^{l}\big)$ with the cosine between output space gradients $\rho_{o}=\cos\!\big(\nabla_{o}\mathcal{L}^{w},\nabla_{o}\mathcal{L}^{l}\big)$ across timesteps and prompts. Steps with near zero norms are filtered to avoid numerical artifacts. We summarize the absolute gap $\Delta\rho_{t}=\big|\rho_{\theta}(t)-\rho_{o}(t)\big|$ and observe small values of $\Delta\rho_{t}$ for the majority of steps, which supports Eq. 20 with $J^{w}{J^{l}}^{\top}\approx I$.
Appendix B Second-Order Considerations of SDPO
It is important to note that our guarantee is based on a first-order (linear) approximation of the loss landscape. In reality, the true change in $\mathcal{L}^{w}$ after an update includes higher-order terms: $\Delta\mathcal{L}^{w}=\nabla_{\theta}{\mathcal{L}^{w}}^{\top}\Delta\theta+\tfrac{1}{2}\Delta\theta^{\top}H^{w}\Delta\theta+\mathcal{O}(|\Delta\theta|^{3})$, where $H^{w}$ is the Hessian of $\mathcal{L}^{w}$. Diffusion-SDPO does not explicitly account for the $\frac{1}{2}\Delta\theta^{\top}H^{w}\Delta\theta$ term. If the curvature (eigenvalues of $H^{w}$) is large or the step $\Delta\theta$ is not infinitesimally small, it is conceivable that $\mathcal{L}^{w}$ could increase slightly even when the first-order term is zero or negative.
To hedge against curvature, we adopt a margin parameter $\mu\!\in\![0,1]$ and set $\lambda_{\text{safe}}\leftarrow(1-\mu)\,\lambda_{\text{safe}}$ (e.g., $\mu=0.1$ gives $0.9\,\lambda_{\text{safe}}$). This adjustment helps ensure $\Delta\mathcal{L}^{w}$ stays negative even when small second-order effects are present. Empirically, we found that the first-order proxy combined with this margin was adequate for stability since the second-order error term remained small relative to the linear term in our experiments.
Another related source of approximation error is the assumption that $J^{w}{J^{l}}^{\top}\approx I$ in Eq. 15. In pathological cases where the model’s parameterization leads to very different Jacobians for $x^{w}_{0}$ and $x^{l}_{0}$, the output-space dot product ${g^{w}}^{\top}g^{l}$ might not perfectly predict the sign of the parameter-space dot product. In practice, however, we expect ${g^{w}}^{\top}g^{l}$ to be a reliable indicator of gradient alignment, since $g^{w}$ and $g^{l}$ live in the same output vector space and capture intuitive per-pixel error correlations between the two samples (see Sec. A for more evidence on this assumption). Our safe update rule can thus be viewed as a principled, efficient heuristic that steers the optimization to avoid obvious “bad” directions that would hurt the preferred outcome. By explicitly safeguarding $\mathcal{L}^{w}$ from increases, Diffusion-SDPO favors a more conservative yet effective alignment, in contrast to methods that risk overshooting in pursuit of preference satisfaction. Generally, our approach leads to a more stable training process that maintains or even improves the quality of winner outputs while still driving down the relative loss of loser outputs, ultimately yielding a well-aligned diffusion model.
Appendix C LLM Usage Disclosure
We used the large language model (LLM) only to aid English writing quality, including grammar correction, style polishing, and minor rephrasing at the sentence or paragraph level. The LLM did not generate research ideas, algorithms, proofs, datasets, code, experimental designs, figures, tables, statistical analyses, or results. All technical content, claims, and conclusions were created and verified by the authors. To reduce risk of factual errors, every LLM-suggested edit was reviewed by at least one author and cross-checked against our methods, experiments, and citations. No undisclosed prompts, hidden instructions, or external links intended to influence the review process were included in the submission. The authors take full responsibility for all content in this paper, including any text edited with LLM assistance, and the LLM is not an author or contributor.
Appendix D Full Experimental Results
We report the full reward score comparisons on Pick-a-Pic V2, HPS V2, and PartiPrompts in Tables 6 and 7. Figure 4 analyzes the sensitivity of $\mu$, the only hyperparameter introduced by our method, by plotting HPS V2 and PickScore as functions of $\mu$. We also provide additional qualitative results in Fig. 5 and 6. Taken together, these quantitative and qualitative results indicate that SDPO consistently improves preference alignment while maintaining, and in many cases enhancing, visual quality.

### v2

@@ 6 Conclusions and Limitations
In this paper, we presented Diffusion-SDPO, a safeguarded preference optimization scheme that stabilizes DPO-style diffusion finetuning by preserving the preferred branch while improving preference matching. The method scales the loser gradient by its alignment with the winner and guarantees, to first order, that the winner’s reconstruction loss does not increase. Across SD 1.5, SDXL (both UNets), and Ovis-U1 (DiT), our method yields consistent improvements on automated preference, aesthetic, and prompt-alignment metrics with negligible computational overhead, while remaining model-agnostic, straightforward to implement, and applicable to multiple DPO variants.
However, the safeguard is derived from a first-order approximation, and its validity weakens when the loss landscape exhibits strong curvature. Moreover, the coefficient $\rho$ (cf. Eq. 16) used to estimate gradient alignment can be noisy or biased, which leads to imperfect scaling of loser gradients. Nevertheless, our empirical results indicate that with a suitably chosen $\mu$, this estimation is accurate enough in practice and the safeguard behaves as intended. Future work includes developing second-order or trust-region safeguards to maintain robust winner preservation under diverse training dynamics.
Appendix A Second-Order Considerations of SDPO
Our theoretical guarantee for Diffusion-SDPO is explicitly first order: it controls the sign of the linear term in the Taylor expansion of the winner loss. In practice, the true change in $\mathcal{L}^{w}$ after an update is
$\Delta\mathcal{L}^{w}=\nabla_{\theta}{\mathcal{L}^{w}}^{\top}\Delta\theta+\tfrac{1}{2}\Delta\theta^{\top}H^{w}\Delta\theta+\mathcal{O}(\|\Delta\theta\|^{3}),$ (18)
where $H^{w}$ is the Hessian of $\mathcal{L}^{w}$ and $\Delta\theta$ is the parameter update induced by the SDPO objective. The analysis in main text ensures that the first-order term $\nabla_{\theta}{\mathcal{L}^{w}}^{\top}\Delta\theta$ is nonpositive under the safe choice of $\lambda$. However, the quadratic term $\tfrac{1}{2}\Delta\theta^{\top}H^{w}\Delta\theta$ can in principle be positive when local curvature is large, so $\mathcal{L}^{w}$ might still increase slightly when the step is not infinitesimal. In other words, SDPO controls the direction of the update with respect to the gradient of $\mathcal{L}^{w}$, while the Hessian governs how quickly the loss can bend back upward along that direction.
The role of the slack parameter $\mu$ to compute $\lambda_{\text{safe}}$ can then be understood as an implicit trust-region style safeguard. Recall that the parameter-space analysis yields an upper bound on $\lambda$, and our output-space scheme uses
$\lambda_{\text{safe}}=\frac{(1-\mu)\|g^{w}\|_{2}^{2}}{g^{w\top}g^{l}},$ (19)
which shrinks the allowable contribution of the loser branch whenever $g^{w\top}g^{l}>0$. Multiplying by $(1-\mu)$ reduces the contribution of the loser gradient in the update. In the regime where $\nabla_{\theta}\mathcal{L}^{w}{}^{\top}\nabla_{\theta}\mathcal{L}^{l}>0$, the first-order change in Eq. 20,
$\Delta\mathcal{L}^{w}\!\!\approx\!\!\nabla_{\theta}{\mathcal{L}^{w}}^{\top}\!\!\Delta\theta\!=\!\!-\eta\Big(\!\|\nabla_{\theta}\mathcal{L}^{w}\|_{2}^{2}\!-\!\lambda\nabla_{\theta}{\mathcal{L}^{w}}^{\top}\!\nabla_{\theta}\mathcal{L}^{l}\!\Big),$ (20)
is monotonically increasing as a function of $\lambda$ (note that both $\lambda$ and $\eta$ are >0>0), so replacing $\lambda$ by $(1-\mu)\lambda$ makes $\Delta\mathcal{L}^{w}$ strictly smaller.
At the same time, we can make the dependence on the loser direction explicit. The SDPO update can be written as
@@ Appendix A Second-Order Considerations of SDPO
$\Delta\theta(\lambda)\!=\!-\eta\big(\nabla_{\theta}\mathcal{L}^{w}\!\!\!-\!\!\lambda\nabla_{\theta}\mathcal{L}^{l}\big)\!=\!\underbrace{-\eta\,\nabla_{\theta}\mathcal{L}^{w}}_{\Delta\theta(0)}\!\!+\!\!\!\!\!\underbrace{\eta\lambda\,\nabla_{\theta}\mathcal{L}^{l}}_{\text{loser-induced part}}\!\!\!\!.$ (21)
The term $\eta\lambda\,\nabla_{\theta}\mathcal{L}^{l}$ is the $\lambda$-dependent increment of the update along the loser direction. Replacing $\lambda$ by $(1-\mu)\lambda$ scales this increment to $\eta(1-\mu)\lambda\,\nabla_{\theta}\mathcal{L}^{l}$, so its norm is multiplied by $(1-\mu)$ while the baseline part $\Delta\theta(0)$ is unchanged. In particular, the portion of the step norm that is directly attributable to the loser branch contracts linearly with $(1-\mu)$. The quadratic term then decomposes as
$\Delta\theta(\lambda)^{\top}H^{w}\Delta\theta(\lambda)=\Delta\theta(0)^{\top}H^{w}\Delta\theta(0)$ (22)
$\!\!\!\!\!\!\quad+2\eta\lambda\,(\nabla_{\theta}\mathcal{L}^{l})^{\top}H^{w}\Delta\theta(0)\!+\!\eta^{2}\lambda^{2}\,(\nabla_{\theta}\mathcal{L}^{l})^{\top}\!H^{w}\nabla_{\theta}\mathcal{L}^{l}\!,$
where the last two terms capture the curvature contribution associated with the loser direction. Under a mild local spectral bound $\|H^{w}\|_{2}\leq\Lambda$, we can control the quadratic term via
$\bigl|v^{\top}H^{w}v\bigr|\leq\|H^{w}\|_{2}\,\|v\|_{2}^{2}\leq\Lambda\,\|v\|_{2}^{2},$ (23)
for any vector $v$. Applying this with $v=\Delta\theta\bigl((1-\mu)\lambda\bigr)$ and using the decomposition in Eq. (21) yields
$\Bigl|\tfrac{1}{2}\Delta\theta\bigl((1-\mu)\lambda\bigr)^{\top}H^{w}\Delta\theta\bigl((1-\mu)\lambda\bigr)\Bigr|$ (24)
$\leq\tfrac{1}{2}\Lambda\bigl\|\Delta\theta\bigl((1-\mu)\lambda\bigr)\bigr\|_{2}^{2}$ (25)
$=\tfrac{1}{2}\Lambda\bigl\|\Delta\theta(0)+\eta(1-\mu)\lambda\,\nabla_{\theta}\mathcal{L}^{l}\bigr\|_{2}^{2}$ (26)
$\leq\tfrac{1}{2}\Lambda\Bigl(\bigl\|\Delta\theta(0)\bigr\|_{2}+\eta(1-\mu)|\lambda|\,\bigl\|\nabla_{\theta}\mathcal{L}^{l}\bigr\|_{2}\Bigr)^{2}$ (27)
$\leq\tfrac{1}{2}\Lambda\Bigl(\bigl\|\Delta\theta(0)\bigr\|_{2}+\eta|\lambda|\,\bigl\|\nabla_{\theta}\mathcal{L}^{l}\bigr\|_{2}\Bigr)^{2}.$ (28)
The last line coincides with the corresponding spectral-norm bound obtained for $\Delta\theta(\lambda)$ before contraction. Since $(1-\mu)|\lambda|\leq|\lambda|$ for $\mu\in[0,1]$, the above inequalities show that the curvature contribution under the contracted update is controlled by an upper bound that is no larger, and typically strictly smaller, than the original one. We therefore claim that SDPO shrinks the worst-case curvature impact in the loser direction, making it less likely that higher-order effects overturn the negative first-order change.
This perspective also clarifies the interaction between SDPO and the base optimizer. In our experiments we use small learning rates and large gradient accumulation, so the effective $\|\Delta\theta\|$ per update is modest even for high-capacity backbones. In this regime, the second-order contribution behaves as a small perturbation around the controlled first-order term. When combined with the slack contraction, this explains why we observe almost monotone or gently decreasing $\mathcal{L}^{w}$ trajectories, rather than the oscillatory behavior that would indicate strong curvature dominating the dynamics.
Empirically, we can also see second-order effects indirectly through the sensitivity of SDPO to $\mu$. If curvature induced large and frequent violations of the first-order approximation, performance would vary sharply with small changes in $\mu$ because the balance between the linear and quadratic terms would be fragile. Instead, we observe broad plateaus where both HPS V2 and PickScore remain near-optimal over wide intervals of $\mu$. This suggests that, in the regions visited during training, the winner loss is reasonably well approximated by its first-order geometry along SDPO updates, and the quadratic term acts as a small correction rather than a dominant force.
Finally, the approximate match between parameter-space and output-space trajectories offers further evidence that higher-order interactions are not pathological in practice. The parameter-space variant implicitly incorporates more of the true curvature structure, yet its behavior can be closely reproduced by the cheaper output-space scheme with a suitably chosen slack. This indicates that most of the relevant geometry is already captured by the alignment between $g^{w}$ and $g^{l}$, and that the remaining second-order discrepancy can be effectively absorbed into a scalar safety margin. A fully second-order SDPO variant that explicitly constrains $\Delta\theta^{\top}H^{w}\Delta\theta$ would be an interesting direction for future work, but our results suggest that the present first-order safeguard, combined with the slack $\mu$, already offers a favorable trade-off between theoretical control, computational cost, and empirical stability.
Table 6: Full reward and structured alignment results for FLUX.1-dev. We report automatic rewards (PickScore, HPS V2, LAION Aesthetics, CLIP, and ImageReward) as well as structured alignment metrics (GenEval and DPG-Bench) for the base model and its preference-aligned variants. The row marked with $\S$ corresponds to Diff.-DPO trained with a $0.01\times$ learning rate.
Method PickScore($\uparrow$) HPS V2($\uparrow$) Aesthetics($\uparrow$) CLIP($\uparrow$) ImageReward($\uparrow$) GenEval($\uparrow$) DPG-Bench($\uparrow$)
FLUX.1-dev 0.2252 0.2889 6.0735 0.3498 0.9916 0.66 85.75
Diff.-DPO 0.1998 0.2467 5.2857 0.2573 -0.7788 0.23 58.97
Diff.-DPO§ 0.2253 0.2881 6.0701 0.3501 0.9938 0.66 85.78
Diff.-DPO + SDPO 0.2314 0.2991 6.1910 0.3566 1.1108 0.71 86.36
Appendix B Extension: SDPO on FLUX.1-dev
Model.
To further test the generality of SDPO, we also apply it to FLUX.1-dev [flux], a 12B-parameter rectified-flow DiT [dit] model designed for text-to-image generation. FLUX.1-dev operates in latent space with a flow-matching objective rather than the noise-prediction objective used in SD 1.5 and SDXL, but it still exposes a time-conditioned image generator whose outputs can be scored by the same preference metrics. In our experiments, we start from the publicly released FLUX.1-dev checkpoint and keep the official tokenizer, text encoder, VAE, and sampling configuration unchanged. We integrate Diffusion-DPO and SDPO at the loss level only, treating the model’s velocity (or denoising) prediction at each time step as the output on which we define the per-step winner and loser residuals.
Datasets.
For DPO training on FLUX.1-dev, we construct an in-house preference dataset of 186K high-quality DPO examples. The images and texts cover a broad range of domains (portraits, everyday objects, indoor and outdoor scenes, concept art, products, etc.), making the corpus suitable for general-purpose alignment rather than a narrow style or category. To improve the textual side of supervision, we apply a vision–language model (Qwen-VL-Max [qwen_vl_max]) to re-caption short, noisy, or underspecified prompts, yielding richer and more semantically faithful descriptions while preserving the original user intent.
Training Details.
We finetune FLUX.1-dev with Diffusion-DPO and with our SDPO-augmented variant using TorchTitan in a hybrid sharded data-parallel (HSDP) configuration. Within each 8-GPU group we apply fully sharded data parallelism (FSDP) to partition model parameters, and we use 2-way data parallelism across groups, resulting in 16 NVIDIA A100 GPUs in total. Training is performed at a resolution of $1024\times 1024$ with a global batch size of 512 and a learning rate of $1\times 10^{-5}$. For the DPO temperature, we set $\beta_{\text{DPO}}=1000$. For SDPO, we choose a slack of $\mu=0.99$, which empirically yields effective safe scaling coefficients $\lambda_{\text{safe}}\approx 0.1$ at typical training steps. Under this configuration, finetuning completes in approximately three days on 16 A100 GPUs.
Evaluation.
We evaluate FLUX.1-dev using the Pick-a-Pic V2 [pap] prompt sets with automatic reward metrics (PickScore [pap], HPS V2 [hpsv2], LAION Aesthetics [laion-aesthetics], CLIP [clip], and ImageReward [imagereward]), and additionally report its structured alignment performance on GenEval [geneval] and DPG-Bench [dpg_bench]. During inference, we adopt a 50-step sampler and follow the FLUX.1-dev model card by using a classifier-free guidance scale of $3.5$ for all variants to ensure a fair comparison.
Results.
Table 6 reveals a clear three-regime behavior on FLUX.1-dev. At the default learning rate ($1\times 10^{-5}$), naive Diffusion-DPO catastrophically degrades the model: all five reward metrics drop, and both GenEval and DPG-Bench scores collapse. This suggests that on a high-capacity rectified-flow backbone, aggressively enlarging the preference margin without any safeguard on the winner branch can drive the parameters far outside the local basin of the pretrained solution.
When we instead reduce the learning rate by a factor of $0.01$ (Diff.-DPO§ with learning rate $1\times 10^{-7}$), the collapse disappears and all metrics revert to nearly the base FLUX.1-dev level, indicating that the optimization has become so conservative that it behaves almost like an identity map and fails to extract meaningful gains from the preference supervision.
Taken together, these two extremes highlight a practical tension when directly applying Diffusion-DPO to large rectified-flow models: in naive settings, one tends to end up either in an aggressive regime that destabilizes the pretrained generator, or in an overly conservative regime where the model remains effectively unchanged.
In contrast, Diff.-DPO + SDPO improves over the base model on all reported metrics while keeping the original, high learning rate. PickScore, HPS V2, Aesthetics, CLIP, and ImageReward all increase, and both GenEval and DPG-Bench scores also rise, indicating that SDPO not only enhances perceptual quality but also strengthens structured text–image alignment.
Appendix C Other Experimental Results
We report the full reward score comparisons on Pick-a-Pic V2, HPS V2, and PartiPrompts in Tables 7 and 8, and summarize the win-rate comparison of SDXL models in Table 9. Across all datasets, adding SDPO on top of Diffusion-DPO, DSPO, or DMPO consistently shifts the metric profile in the same direction as in the main text: PickScore, HPS V2, and ImageReward improve while aesthetic scores are preserved or slightly enhanced, indicating that SDPO strengthens preference alignment without paying a quality penalty.
Additional qualitative examples in Fig. 7 and Fig. 8 further illustrate this effect on both UNet and DiT backbones, showing sharper details, more faithful layouts, and fewer artifacts compared to their non-SDPO counterparts. Finally, Fig. 9 visualizes the temporal dynamics: as training progresses, baseline Diffusion-DPO drifts toward over-saturated or distorted images, illustrating how the winner loss can keep increasing even while the winner-loser margin becomes more negative when optimization focuses purely on enlarging the margin. In contrast, SDPO maintains coherent structure and style across steps, thanks to its winner-preserving update rule and the stabilizing effect of the safe scaling $\lambda_{\text{safe}}$. Taken together, these quantitative and qualitative results support our central claim that SDPO acts as a plug-in, geometry-aware safeguard that consistently improves preference metrics while maintaining, and in many cases enhancing, the visual quality of diffusion generations.